Dr. Stallter has been teaching basic statistics for many years, she knows that 80 percent of the students will complete the assigned problems. She also has determined that among those who completed their assignments, 90 percent will pass the course. Among those students who do not do their homework, 60 percent will pass. Mike Fishbaugh took statistics last semester from Dr. Stallter and received a passing grade. What is the probability that he completed the assignments?
1
Expert's answer
2012-06-15T11:04:20-0400
People who did homeworks and passed the course: 0.8*0.9=0.72
People who did not homeworks and passed the course: 0.2*0.6=0.12
Probability, that Mike did his home works is: 0.72/(0.12+0.72)=0.857